OpenStudy (anonymous):

what is the average velocity of an arrow that travels 12m in 0.15s

OpenStudy (rane):

v=s/t

OpenStudy (anonymous):

can u answer it plz

OpenStudy (rane):

ok since v=dispalacement/ time = d/t u r already given s and t, u just need to subsitude the both value in the equation, it will bw: v= 12/0.5

OpenStudy (anonymous):

thx so much

OpenStudy (rane):

you r welcome

OpenStudy (anonymous):

what about 'if a conoe has an average velocity of 4.2m/s what is the displacement after 12s

OpenStudy (rane):

same thing. v=s/t 4.2=s/12...so u have to rearrange it, take the s on the other side and bring 4.2 where s was: s= vt = 4.2*12

OpenStudy (anonymous):

where did u get 12 from

OpenStudy (anonymous):

my mad it was 25 not 12

OpenStudy (anonymous):

bad*

OpenStudy (rane):

so u then just replace 12 by 25 its still the time.